OpenStudy (anonymous):

The container that holds the water for the football team is 2/5 full. After pouring in 8 gallons of water, it is 2/3 full. How many gallons can the container hold?

OpenStudy (flashp):

2/5 of 30 = 60/5 = 12. Add 8 gallons to 12 and we have 20. 2/3 of 30 = 60/3 = 20 which means you answer would be 30

OpenStudy (alekos):

But how did you get 30 in the first place? Just a wild guess?

OpenStudy (alekos):

Using algebra 2x/5 + 8 = 2x/3 where x is the volume of the container Solve for x